OpenStudy (anonymous):

Help Trigonometric substitution \[\int\limits_{5}^{10}\sqrt{100-x ^{2}}dx\]

hartnn (hartnn):

same hint : a^2-x^2 put x= a sin theta

OpenStudy (anonymous):

let x = 10sin(x)

OpenStudy (anonymous):

\[\sqrt{100-(10\sin \theta)^{2}}dx\]

OpenStudy (anonymous):

\[\sqrt{100-100\sin ^{2}x}\]

OpenStudy (anonymous):

Then use a trig identity

OpenStudy (anonymous):

then dx= 10 cos t

OpenStudy (anonymous):

Yes, let me work out the full problem real quick

hartnn (hartnn):

yes, ans what does sqrt (....) simplify to ?

OpenStudy (anonymous):

so will it eventually just be 100 cos^2 t

hartnn (hartnn):

yes. and you know the cos 2x formula ?

OpenStudy (anonymous):

cos^2 t = (1+cos2 t)/2 that one?

OpenStudy (anonymous):

yes

hartnn (hartnn):

yes, that should make it easier to solve now :)

OpenStudy (anonymous):

alright thanks so much again

hartnn (hartnn):

ask if you get stuck again anywhere....